Grade 12Physical Chemistry

2H2O––––2H2O+O2
A)rate of formation of H2O
B)rate of disappearance of H2O2

Surjeet Singh Kushwah
8 Years agoGrade 12

1 Answer

Rituraj Tiwari
5 Years ago

Let's break down the reaction you've presented: 2H2O2 → 2H2O + O2. This equation indicates the decomposition of hydrogen peroxide (H2O2) into water (H2O) and oxygen gas (O2). To address your question regarding the rates of formation and disappearance in this reaction, we’ll look at the stoichiometry involved.

Understanding Reaction Rates

In chemical reactions, the rate at which reactants disappear and products form can be described in terms of their stoichiometric coefficients. For the given reaction, we can identify how the rates relate to each other based on the coefficients in front of each substance.

Rate of Formation of H2O

The rate of formation of water (H2O) is directly tied to the stoichiometry of the reaction. According to the balanced equation, for every 2 moles of hydrogen peroxide that decompose, 2 moles of water are produced. Therefore, we can express the rate of formation of H2O as:

  • Rate of formation of H2O = (1/2) × Rate of disappearance of H2O2

This means that the rate at which water is formed will be equal to the rate at which hydrogen peroxide is consumed because they have the same stoichiometric coefficient (2). Specifically, if the hydrogen peroxide disappears at a rate of x, water is formed at the same rate x.

Rate of Disappearance of H2O2

Now, let's consider the rate of disappearance of hydrogen peroxide (H2O2). The stoichiometry shows that for every 2 moles of H2O2 that react, 1 mole of O2 is produced and 2 moles of H2O are formed. Thus, the rate of disappearance of H2O2 can be expressed as:

  • Rate of disappearance of H2O2 = (1/2) × Rate of formation of O2

This indicates that the rate of hydrogen peroxide disappearing is double the rate of oxygen gas formation, given that 2 moles of H2O2 produce only 1 mole of O2. So, if we express the rates mathematically, we can summarize the relationships as follows:

  • Rate of formation of H2O = Rate of disappearance of H2O2
  • Rate of formation of O2 = (1/2) × Rate of disappearance of H2O2

Putting It All Together

In conclusion, the relationships among the rates of the reactants and products are dictated by the coefficients in the balanced chemical equation. If you know the rate of one component, you can easily calculate the rates of the others using the stoichiometric relationships. This understanding is crucial for applications in kinetics, where knowing how fast reactants turn into products can dictate the conditions needed for a reaction to be efficient or safe.